OpenStudy (anonymous):

What is the equation of the line which includes points (1, 5) and (-2, -7)?

OpenStudy (anonymous):

hi :)

OpenStudy (anonymous):

helloo

OpenStudy (anonymous):

okay so so first you should find the slope of the line [slope= (y2-y1)/(x2-x1)]

OpenStudy (anonymous):

try it and tell me what you get :)

OpenStudy (anonymous):

how do i find the slopee

OpenStudy (anonymous):

[slope= (y2-y1)/(x2-x1)]

OpenStudy (anonymous):

??? how did you get that

OpenStudy (anonymous):

\[y _{2} - y _{1}\over x _{2}-x _{1}\]

OpenStudy (anonymous):

thats the equation to find slope :)

OpenStudy (anonymous):

okaii then what

OpenStudy (anonymous):

so now you put in the numbers...\[-7-5\over-2-1\]

OpenStudy (anonymous):

so i would do -7-5/-2-1

OpenStudy (anonymous):

so you combine the terms and get...\[-12\over-3\]

OpenStudy (anonymous):

now that can be simplified even more....\[-4\over-1\]

OpenStudy (anonymous):

and the negatives get canceled out and the one goes away \[4\]

OpenStudy (anonymous):

so now you have your slope :)

OpenStudy (anonymous):

im confused because thats not what the question asks

OpenStudy (anonymous):

yes but first you need to find the slope to write the equation of the line :)

OpenStudy (anonymous):

okay so the equation you need to put it into is \[y=mx+b\]

OpenStudy (anonymous):

now you just found m (the slope)

OpenStudy (anonymous):

so now choose either one of the two points you started off with :)

OpenStudy (anonymous):

(1,5) or (-2,-7) ? :) which one do you want to use??

OpenStudy (anonymous):

Let x and y be on the line. Equate the calculation for the slope using the first point and the second point respectively.\[\frac{y-5}{x-1}\text{=}\frac{y+7}{x+2} \]Solve the above for y.\[y=4 x+1 \]

OpenStudy (anonymous):

you just...gave..the answer..........

OpenStudy (anonymous):

and the method.

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

thank yu :) cause i didnt understand a thing he was saying yu made it simple and easy

OpenStudy (anonymous):

You're welcome.